OpenStudy (anonymous):

Can someone please tell me what they get to this problem Find the derivative of the function j(x)=square root x^10-x^8

OpenStudy (inkyvoyd):

If you say what you got first :)

OpenStudy (anonymous):

ok I got\[\frac{ 5x^9-4x^7 }{ (x ^{10}-x^8)\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

that is supposed to be raised to 3/2 sorry

OpenStudy (anonymous):

is that what you got?

OpenStudy (amistre64):

\[D_u[\sqrt{u}]=\frac{u'}{2\sqrt{u}}\]

OpenStudy (anonymous):

you mean the answer should be \[\frac{ 1 }{ 2\sqrt{10x^9-8x^7} }\]

OpenStudy (amistre64):

no

OpenStudy (amistre64):

your answer was almost correct, the 3/2 should have been 1/2 to indicate a sqrt

OpenStudy (anonymous):

ok your saing that the answer I posted previously was correct but the denominator needs to be square rooted

OpenStudy (amistre64):

yes, .. or another way to write a sqrt is with an exponent of 1/2 .... not 3/2

OpenStudy (anonymous):

ok gotcha thanks

OpenStudy (amistre64):

youre welcome